# Effect of the scalar unparticle and polarization at muon colliders through the exclusive W boson hadronic decays in the final state in the Randall-Sundrum model

### v3

@@ I Introduction
      v) We evaluate the benchmark background $(\Lambda_{U},d_{U})$ in Fig.5. The parameters are chosen as above, $P_{\mu^{-}}=0.8,P_{\mu^{+}}=-0.8$, $\sqrt{s}=13$ TeV, $\Delta k_{\gamma}=0.190$, $\lambda_{\gamma}=0.061$, $\Delta k_{Z}=0.093$, $\lambda_{Z}=0.065$. From the figure, we can see that with $(\Lambda_{U},d_{U})$ $=(1TeV,1.9)$, the cross-sections in final states reach the maximum values. This result is similar to Ref.[25].
      vi) For background, we choose the parameters as $P_{\mu^{-}}=0.8,P_{\mu^{+}}=-0.8$, $\sqrt{s}=13$ TeV, $\Delta k_{\gamma}=0.190$, $\lambda_{\gamma}=0.061$, $\Delta k_{Z}=0.093$, $\lambda_{Z}=0.065$, $\Lambda_{U}=1$ TeV, $d_{U}=1.1$. Using the integrated luminosity scaling of a high energy muon collider 140 $fb^{-1}$ [40], some values $S/\sqrt{S+B}$ are given in Table.1. The results show that the pair production of mesons and photons can be detected with a small significance level. Therefore, we evaluate the integrated luminosity value correspond to $1\sigma,3\sigma$ for the final states included scalar unparticle propagator in detail in Table.2. The cross-sections in the propagators are considered in detail in Table.3, which shows that the contribution of t-channel (neutrino propagator) is larger than that of s-channel under the same conditions. In case $\mu^{-}$ beam is left polarized, $\mu^{+}$ beam is right polarized and vice versa, there in only the s-channel contribution. The forward-backward asymmetry $A_{FB}$ values in case of the different polarization coefficients of $\mu^{-},\mu^{+}$ beams are indicated in Table.4. At fixed collision energy and in case of $(P_{\mu^{-}},P_{\mu^{+}})$ $=(1,-1),(0.8,-0.8),(0.6,-0.6),(1,1)$, the forward-backward asymmetry $A_{FB}$ is about 0.0050076, 0.000255, 0.000254, 0.000252, respectively.
Table 1: $S/\sqrt{S+B}$ at 13 TeV muon collider with 140 $fb^{-1}$ integrated luminosity.
Process $(P_{\mu^{-}},P_{\mu^{+}})$ $\sigma_{B}$ (fb) $\sigma_{S}$ (fb) $S/\sqrt{S+B}$
(1, -1) 3.02273 ×$\times 10^{-13}$ 2.8179 ×$\times 10^{-7}$ 0.0062
$\mu^{+}\mu^{-}\rightarrow W^{+}W^{-}\rightarrow\pi^{-}\pi^{+}\gamma\gamma$ (0.8, -0.8) 0.00004 0.00029 0.1888
(0.6, -0.6) 0.00007 0.00052 0.2533
(1, 1) 0.00022 0.00164 0.4499
(1, -1) 2.41986 ×$\times 10^{-13}$ 2.25589 ×$\times 10^{-7}$ 0.0056
$\mu^{+}\mu^{-}\rightarrow W^{+}W^{-}\rightarrow K^{-}K^{+}\gamma\gamma$ (0.8, -0.8) 0.00003 0.00023 0.1687
(0.6, -0.6) 0.00005 0.00042 0.2292
(1, 1) 0.00018 0.00131 0.4015
(1, -1) 2.26412 ×$\times 10^{-12}$ 2.1107 ×$\times 10^{-6}$ 0.0172
$\mu^{+}\mu^{-}\rightarrow W^{+}W^{-}\rightarrow\rho^{-}\rho^{+}\gamma\gamma$ (0.8, -0.8) 0.00031 0.00221 0.5209
(0.6, -0.6) 0.00054 0.00392 0.6945
(1, 1) 0.00169 0.01226 1.2282
Table 2: The integrated luminosity value $\mathcal{L}$ correspond to $1\sigma,3\sigma$ for the final states included scalar unparticle propagator .
Process $(P_{\mu^{-}},P_{\mu^{+}})$ $1\sigma$ $3\sigma$
(1, -1) 3.54 ×$\times 10^{6}$ $fb^{-1}$ 3.19 ×$\times 10^{7}$ $fb^{-1}$
$\mu^{+}\mu^{-}\rightarrow W^{+}W^{-}\rightarrow\pi^{-}\pi^{+}\gamma\gamma$ (0.8, -0.8) 3923 $fb^{-1}$ 35315 $fb^{-1}$
(0.6, -0.6) 2182 $fb^{-1}$ 19637 $fb^{-1}$
(1, 1) 691 $fb^{-1}$ 6224 $fb^{-1}$
(1, -1) 4.43 ×$\times 10^{6}$ $fb^{-1}$ 3.99 ×$\times 10^{7}$ $fb^{-1}$
$\mu^{+}\mu^{-}\rightarrow W^{+}W^{-}\rightarrow K^{-}K^{+}\gamma\gamma$ (0.8, -0.8) 4914 $fb^{-1}$ 44234 $fb^{-1}$
(0.6, -0.6) 2664 $fb^{-1}$ 23979 $fb^{-1}$
(1, 1) 868 $fb^{-1}$ 7814 $fb^{-1}$
(1, -1) 473777 $fb^{-1}$ 4.26 ×$\times 10^{6}$ $fb^{-1}$
$\mu^{+}\mu^{-}\rightarrow W^{+}W^{-}\rightarrow\rho^{-}\rho^{+}\gamma\gamma$ (0.8, -0.8) 516 $fb^{-1}$ 4643 $fb^{-1}$
[…]
(1, 1) 93 $fb^{-1}$ 835 $fb^{-1}$
Table 3: Cross-sections in the propagators at 13 TeV muon collider .
Process $(P_{\mu^{-}},P_{\mu^{+}})$ $\sigma_{\phi,h,U}$ (fb) $\sigma_{\gamma,Z}$ (fb) $\sigma_{\nu}$ (fb)
(1, -1) 2.81791 ×$\times 10^{-7}$ 0 0
$\mu^{+}\mu^{-}\rightarrow W^{+}W^{-}\rightarrow\pi^{-}\pi^{+}\gamma\gamma$ (0.8, -0.8) 2.29137 ×$\times 10^{-7}$ 0.000031 0.000136
(0.6, -0.6) 1.84141 ×$\times 10^{-7}$ 0.000055 0.000242
(1, 1) 0 0.000172 0.000758
(1, -1) 2.25589 ×$\times 10^{-7}$ 0 0
$\mu^{+}\mu^{-}\rightarrow W^{+}W^{-}\rightarrow K^{-}K^{+}\gamma\gamma$ (0.8, -0.8) 1.83436 ×$\times 10^{-7}$ 0.000025 0.000109
(0.6, -0.6) 1.47415 ×$\times 10^{-7}$ 0.000044 0.000194
(1, 1) 0 0.000139 0.000607
(1, -1) 2.1107 ×$\times 10^{-6}$ 0 0
$\mu^{+}\mu^{-}\rightarrow W^{+}W^{-}\rightarrow\rho^{-}\rho^{+}\gamma\gamma$ (0.8, -0.8) 1.7163 ×$\times 10^{-6}$ 0.000235 0.001022
(0.6, -0.6) 1.37928 ×$\times 10^{-6}$ 0.000418 0.001817
(1, 1) 0 0.001307 0.005679
Table 4: Forward-backward asymmetry $A_{FB}$ in case of the different polarization coefficients of $\mu^{-},\mu^{+}$ beams.
$\sqrt{s}$ 10 TeV 13 TeV 14TeV
$A_{FB}$ ($(P_{\mu^{-}},P_{\mu^{+}})$ =(1, -1)) 0.005076 0.005076 0.005076
$A_{FB}$ ($(P_{\mu^{-}},P_{\mu^{+}})$ = (0.8, -0.8)) 0.000255 0.000255 0.000255
$A_{FB}$ ($(P_{\mu^{-}},P_{\mu^{+}})$ = (0.6, -0.6)) 0.000254 0.000254 0.000254
$A_{FB}$ ($(P_{\mu^{-}},P_{\mu^{+}})$ = (1,1)) 0.000252 0.000252 0.000252
III Conclusion
@@
In this paper, by using Feynman diagram techniques we have evaluated the effect of the scalar unparticle and polarization at the muon colliders through the exclusive W boson hadronic decays in the RS model. The results show that with fixed collision energies, the total cross-sections for hadronic productions in final states depend strongly on the parameters of the unparticle physics and muon beam polarizes. The cross-sections achieve the maximum value when both of muon beams polarize left or right and the minimum value when the $\mu^{-}$ beam polarizes left, the $\mu^{+}$ beam polarizes right and vice versa. In case of the different polarization, the cross section increases as the collision energy increases and it changes insignificantly when the $\mu^{-}$ beam polarizes left, the $\mu^{+}$ beam polarizes right. With the benchmark background $(\Lambda_{U},d_{U})$ =(1=(1= ( 1TeV,1.9),1.9), 1.9 ), the cross-sections reach the maximum value. In our numerical estimation, the measured forward-backward asymmetry is quiet small which shows the direction to collect the final state from the experiment.
@@
          Finally, we emphasize that in this work we have only considered on a theoretical basis, other problems concerning experiments of exclusive boson hadronic decays , the reader can see in detail in Refs.[40, 42]. Note that these rare processes provide a test bench for the quantum chromodynamics factorization formalism used to calculate cross-sections at high energy colliders (the center-of-mass energy up to tens of TeV), as well as a probe of W boson coupling to quarks and a new way to measure the W boson mass through fully reconstructed decay products.
@@
APPENDIX A: Feynman diagrams for the considered process
Figure 6: Feynman diagrams for $\mu^{+}\mu^{-}\rightarrow W^{+}W^{-}\rightarrow\pi^{-}\pi^{+}\gamma\gamma/K^{-}K^{+}\gamma\gamma/\rho^{-}\rho^{+}\gamma\gamma$ collisions. $M^{\pm}$ stands for the $\pi^{\pm},K^{\pm},\rho^{\pm}$

### v5

@@ II Theoretical framework of the Higgs - radion mixing in Randall-Sundrum model
    i) We evaluate the benchmark parameters $(\Lambda_{U},d_{U})$ in Fig.2. The parameters are taken to be $P_{\mu^{-}}=0.8,P_{\mu^{+}}=-0.8$, $\sqrt{s}=10$ TeV, $\Delta k_{\gamma}=0.190$, $\lambda_{\gamma}=0.061$, $\Delta k_{Z}=-0.061$, $\lambda_{Z}=-0.062$. From the figure, we can see that the cross-sections reach their maximum values at the benchmark signal point $(\Lambda_{U},d_{U})$ $(1\text{TeV},1.9)$. This result is similar to Ref.[32].
Refer to caption Refer to caption Refer to caption
(a) (b) (c)
Figure 2: The total cross-section depends on the ($\Lambda_{U},d_{U}$) in the reactions of (a) $\mu^{+}\mu^{-}\rightarrow W^{+}W^{-}\rightarrow\pi^{+}\pi^{-}\gamma\gamma$, (b) $\mu^{+}\mu^{-}\rightarrow W^{+}W^{-}\rightarrow K^{+}K^{-}\gamma\gamma$ and (c) $\mu^{+}\mu^{-}\rightarrow W^{+}W^{-}\rightarrow\rho^{+}\rho^{-}\gamma\gamma$, respectively. The parameters are chosen as $\sqrt{s}=10$ TeV, $(P_{\mu^{-}},P_{\mu^{+}})=(0.8,-0.8)$, $(\Delta k_{\gamma},\lambda_{\gamma})=(0.190,0.061)$, $(\Delta k_{Z},\lambda_{Z})=(-0.061,-0.062)$.
ii) With the benchmark signal point $(\Lambda_{U},d_{U})=(1\text{TeV},1.9)$, the total cross-section depends on ($\Delta k_{\gamma}$, $\lambda_{\gamma}$) shown in the Fig. 3. The other parameters are chosen as in Fig.2. The figure shows that cross-sections are the largest in the yellow region, the typical values are given by (a) $0.8\times 10^{-9}$ fb for the decay $W^{\pm}\rightarrow\pi^{\pm}\gamma$, (b) $0.53\times 10^{-11}$ fb for the decay $W^{\pm}\rightarrow K^{\pm}\gamma$ and (c) $3.8\times 10^{-9}$ fb for the decay $W^{\pm}\rightarrow\rho^{\pm}\gamma$, respectively.
Refer to caption Refer to caption Refer to caption
(a) (b) (c)
Figure 3: The total cross-section depends on the ($\Delta k_{\gamma},\lambda_{\gamma}$). The benchmark signal point are taken to be $(\Lambda_{U},d_{U})$ $(1\text{TeV},1.9)$. The other parameters are chosen as in Fig.2.
iii) Similar to the Fig.3, the total cross-sections depend on ($\Delta k_{Z}$, $\lambda_{Z}$) are shown in the Fig.4. The typical values of the cross -section are given by (a) $0.810\times 10^{-9}$ fb, (b) $0.535\times 10^{-11}$ fb and (c) $3.870\times 10^{-9}$ fb, respectively.
Refer to caption Refer to caption Refer to caption
(a) (b) (c)
Figure 4: The total cross-section depends on the ($\Delta k_{Z},\lambda_{Z}$). The benchmark signal point are taken to be $(\Lambda_{U},d_{U})$ $(1\text{TeV},1.9)$. The parameters are chosen as in Fig.2.
iv) In Fig.5, the cross-sections are plotted as the function of polarization coefficients ($P_{\mu^{-}},P_{\mu^{+}}$). The parameters are chosen as above, i.e $d_{U}=1.9$, $\Lambda_{U}=1$ TeV, $\sqrt{s}=10$ TeV, $\Delta k_{\gamma}=0.190$, $\lambda_{\gamma}=0.061$, $\Delta k_{Z}=-0.061$, $\lambda_{Z}=-0.062$. The figures indicate that the cross-sections achieve the maximum values when $P_{\mu^{-}}=P_{\mu^{+}}=1$ or 1-1 and the minimum values when $P_{\mu^{-}}=1,P_{\mu^{+}}=-1$ or $P_{\mu^{-}}=-1,P_{\mu^{+}}=1$, respectively.
Refer to caption Refer to caption Refer to caption
(a) (b) (c)
Figure 5: The total cross-section as a function of the polarization coefficients ($P_{\mu^{-}},P_{\mu^{+}}$). The parameters are chosen as $\sqrt{s}=10$ TeV, $(\Lambda_{U},d_{U})$ $(1\text{TeV},1.9)$, $(\Delta k_{\gamma},\lambda_{\gamma})=(0.190,0.061)$, $(\Delta k_{Z},\lambda_{Z})=(-0.061,-0.062)$.
v) The cross-section dependence on the energy is shown in Fig.6. The parameters are chosen as in Fig.5 and the polarization coefficients $(P_{\mu^{-}},P_{\mu^{+}})$ are similar to Ref.[49], i.e $(P_{\mu^{-}},P_{\mu^{+}})$ =(1,1),(0.8,0.8),(0.6,0.6),(1,1)=(1,-1),(0.8,-0.8),(0.6,-0.6),(1,1), respectively. From the figure we can see that the cross-sections increase when the collision energy increases. In case of $\mu^{-}$ beam is left polarized, $\mu^{+}$ beam is right polarized and vice versa, the cross-sections change insignificantly.
Refer to caption Refer to caption Refer to caption
(a) (b) (c)
Figure 6: The total cross-section depends on the collision energy. The parameters are chosen as in Fig.5.
vi) Finally, for more details, we calculate the statistical significance for the exclusive W boson hadronic decays at muon colliders which is defined by [57, 58] $SS=S/\sqrt{S+B}$, where B, S are the number of events for the signal in SM framework and in RS framework including unparticle, anomalous couplings, respectively. The parameters are chosen as in Fig.5 and the polarization coefficients of $(P_{\mu^{-}},P_{\mu^{+}})=(1,1)$ are chosen. Using the integrated luminosity scaling of 10 $\text{ab}^{-1}$ (assuming a 5-year run) [40, 41], the some typical values of the statistical significance are given in Table.1. The analysis reveals that the statistical significance of the exclusive hadronic W boson channel at muon colliders remains small, yielding $\text{SS}=(0.0048\pm 0.00096)\sigma$, $(0.00039\pm 0.00008)\sigma$, and $(0.0105\pm 0.00229)\sigma$ for the decays of $W^{\pm}\rightarrow\pi^{\pm}\gamma$, $W^{\pm}\rightarrow K^{\pm}\gamma$ and $W^{\pm}\rightarrow\rho^{\pm}\gamma$, respectively. However, with the contribution from the RS model including unparticle, and anomalous couplings, the cross-sections are greatly enhanced (the readers can see more details in Table.2). It is worth that there is not universal agreement between these predictions. For example, the statistical significance SS can reach $3\sigma-8\sigma$ in case of using the W branching ratio at the experimental upper bounds of $\mathcal{O}(10^{-6})$ in Ref.[48].
Table 1: Some typical values of the statistical significance for the exclusive W boson hadronic decays at high energy muon colliders with the integrated luminosity of 10 $\text{ab}^{-1}$. The parameters are chosen as in Fig.5
and the polarization coefficients of $(P_{\mu^{-}},P_{\mu^{+}})=(1,1)$ are chosen.
Process $\sigma_{B}$ (fb) $\sigma_{S}$ (fb) $SS(\sigma)$
$\mu^{+}\mu^{-}\rightarrow W^{+}W^{-}\rightarrow\pi^{-}\pi^{+}\gamma\gamma$ (3.823 $\pm$ 0.153 ) $\times 10^{-9}$ (4.354 $\pm 0.174$) $\times 10^{-9}$ 0.0048 $\pm$ 0.00096
$\mu^{+}\mu^{-}\rightarrow W^{+}W^{-}\rightarrow K^{-}K^{+}\gamma\gamma$ (2.524 $\pm$ 0.114) $\times 10^{-11}$ (2.874 $\pm 0.129$) $\times 10^{-11}$ 0.00039 $\pm$ 0.00008
$\mu^{+}\mu^{-}\rightarrow W^{+}W^{-}\rightarrow\rho^{-}\rho^{+}\gamma\gamma$ (18.252 $\pm$ 0.872 ) $\times 10^{-9}$ (20.785 $\pm 0.993$) $\times 10^{-9}$ 0.0105 $\pm$ 0.00229
Table 2: Some typical values in the layers: SM, RS including unparticle and anomalous couplings. The parameters are taken to be $(\Lambda_{U},d_{U})$ $(1\text{TeV},1.9)$ and the polarization coefficients of $(P_{\mu^{-}},P_{\mu^{+}})=(1,1)$ are chosen.
Final states $\sqrt{s}$ $\sigma_{B}(\sigma_{SM})$ (fb) $\sigma_{S}(\sigma_{\text{RS + U + anomalous}})$ (fb)
$\pi^{-}\pi^{+}\gamma\gamma$ 3 (3.450 $\pm$ 0.138 ) $\times 10^{-10}$ (3.929 $\pm 0.157$) $\times 10^{-10}$
10 (3.823 $\pm$ 0.153 ) $\times 10^{-9}$ (4.354 $\pm 0.174$) $\times 10^{-9}$
14 (7.526 $\pm$ 0.301 ) $\times 10^{-9}$ (8.566 $\pm 0.343$) $\times 10^{-9}$
$K^{-}K^{+}\gamma\gamma$ 3 (2.278 $\pm$ 0.103 ) $\times 10^{-12}$ (2.594 $\pm 0.117$) $\times 10^{-12}$
10 (2.524 $\pm$ 0.114) $\times 10^{-11}$ (2.874 $\pm 0.129$) $\times 10^{-11}$
14 (4.968 $\pm$ 0.224 ) $\times 10^{-11}$ (5.655 $\pm 0.255$) $\times 10^{-11}$
$\rho^{-}\rho^{+}\gamma\gamma$ 3 (16.471 $\pm$ 0.787 ) $\times 10^{-10}$ (18.757 $\pm 0.895$) $\times 10^{-10}$
10 (18.252 $\pm$ 0.872 ) $\times 10^{-9}$ (20.785 $\pm 0.993$) $\times 10^{-9}$
14 (35.929 $\pm$ 1.716 ) $\times 10^{-9}$ (40.894 $\pm 1.953$) $\times 10^{-9}$
IV Conclusion
@@
In this paper, by using Feynman diagram techniques we have evaluated the influence of the scalar unparticle and polarization on the exclusive W boson hadronic decays of $W^{\pm}\rightarrow\pi^{\pm}\gamma$, $W^{\pm}\rightarrow K^{\pm}\gamma$ and $W^{\pm}\rightarrow\rho^{\pm}\gamma$ at the high energy muon colliders in the RS model. The results show that with fixed collision energies, the total cross-sections for hadronic productions in final state depend strongly on the parameters of the unparticle physics and muon beam polarizes. With a center-of-mass energy of 10 TeV, the total cross-sections achieve the maximum value when the benchmark signal point as $(\Lambda_{U},d_{U})$ $(1\text{TeV},1.9)$ and the polarization coefficients as $(P_{\mu^{-}},P_{\mu^{+}})=(1,1)$. For more details we have evaluated the statistical significance for the considered process. The numerical values indicate that the effect is greatly enhanced in the future experiments with moderately high energy colliders.
@@
     Finally, we emphasize that in this work we have only considered on a theoretical basis, other problems concerning experiments of exclusive boson hadronic decays, the reader can see in detail in Refs.[48, 50]. It is worth noting that these rare processes provide a test bench for the quantum chromodynamics factorization formalism used to calculate cross-sections at high energy colliders, as well as a probe of W boson coupling to quarks and a new way to measure the W boson mass through fully reconstructed decay products. We will continue to investigate processes resulting in the decay channels of W boson to other meson final states in our subsequent studies.